An enzyme catalyzes conversion of sucrose to fructose. The half-life of the reaction depends on the concentration of sucrose. With an initial concentration of 3.0 M sucrose, t1/2 = 1.36x10–4 s. With an initial concentration of 1.0 M sucrose, t1/2 = 4.5x10–5 s.

a. What is the reaction order in sucrose?

b. Calculate the rate constant for the reaction.
Answer #1

Step (1) using the formula

HALF-LIFE DETERMINATION METHOD In general the half-life of a reaction in which the concentration of all reactants are identic putting the all value in formula

n = [log{(1.36×10-4)/(4.5×10-5)}/log(3.0/1.0)] + 1

n = [log(3)/log(3)] + 1

n = 1+1

n = 2

So

Order of reaction is 2

(b)

t(1/2) = 1/k×an-1

k = 1/t(1/2)×an-1

putting the value in formula

k = 1/(1.36×10-4s)(3.0M)2-1

k = 1/4.08×10-4Ms

k = 2.450×10-3M-1s-1